### v1

# HIF: The hypergraph interchange format for higher-order networks

Abstract
Many empirical systems contain complex interactions of arbitrary size, representing, for example, chemical reactions, social groups, co-authorship relationships, and ecological dependencies. These interactions are known as higher-order interactions and the collection of these interactions comprise a higher-order network, or hypergraph. Hypergraphs have established themselves as a popular and versatile mathematical representation of such systems and a number of software packages written in various programming languages have been designed to analyze these networks. However, the ecosystem of higher-order network analysis software is fragmented due to specialization of each software’s programming interface and compatible data representations. To enable seamless data exchange between higher-order network analysis software packages, we introduce the Hypergraph Interchange Format (HIF), a standardized format for storing higher-order network data. HIF supports multiple types of higher-order networks, including undirected hypergraphs, directed hypergraphs, and simplicial complexes, while actively exploring extensions to represent multiplex hypergraphs, temporal hypergraphs, and ordered hypergraphs. To accommodate the wide variety of metadata used in different contexts, HIF also includes support for attributes associated with nodes, edges, and incidences. This initiative is a collaborative effort involving authors, maintainers, and contributors from prominent hypergraph software packages. This project introduces a JSON schema with corresponding documentation and unit tests, example HIF-compliant datasets, and tutorials demonstrating the use of HIF with several popular higher-order network analysis software packages.
complex network data; higher-order network science; hypergraph; directed hypergraph; simplicial complex; dataset standards
I Background
The growth of network science as a field has been fueled by the availability of large-scale datasets from a variety of sources including the internet [1], co-authorship networks [2, 3], global-scale transportation [4], social interactions [5, 6, 7], biological processes [8, 9], social media [10, 11, 12], and brain connectivity [13], among others. As the plethora of data has grown, standardized formats for pairwise network datasets have been developed such as GraphML [14] and Pajek [15] as well as network dataset repositories such as SNAP [16], Netzschleuder [17], ICON [18], and Konect [19]. These advances, however, have been developed with pairwise networks primarily in mind, which solely model dyadic interactions.
Higher-order networks generalize the notion of pairwise networks, representing interactions of arbitrary size. These networks can more naturally model certain empirical systems such as co-authorship networks, chemical reactions, human social networks, and biological interactions, which abound in multi-way interactions. This can be useful for studying nonlinear dynamical processes [20], the evolution of groups [21], and protein-protein interactions [22], for example. There are many robust software packages for storing, analyzing, and visualizing higher-order networks [23, 24, 25, 26, 27, 28, 29, 30, 31], but no corresponding data standards for facilitating cross-package compatibility. In this paper, co-authored by leaders and representatives of a number of these software packages, we explain the need for such a standard and the context in which it fits; introduce the Hypergraph Interchange Format (HIF) as a data standard for higher-order networks; describe integration with existing higher-order network software libraries; and lastly, demonstrate how it can be used to unlock the strengths of different higher-order network software packages.
I.1 Existing higher-order network data
Higher-order network data remains relatively scattered, with multiple versions of the same datasets often distributed across several repositories and lacking a consistent format. For example, the webpage curated by Austin Benson [32] features many datasets, primarily from the computer science domain, but is no longer actively maintained. There exist several actively maintained data repositories: XGI-DATA [31], which hosts many datasets on Zenodo  [^1] See https://zenodo.org/communities/xgi. and is integrated with the XGI software API; HypergraphRepository [34], which offers a collection of empirical hypergraph datasets along with a web application for filtering and downloading them; and the Hypergraphx [26] software library, which is accompanied by a diverse collection of datasets spanning multiple domains and hypergraph types [35]. Additionally, higher-order network datasets can sometimes be found in traditional network repositories such as SNAP [16], ICON [18], and Netzschleuder [17], typically represented as bipartite graphs. Notably, there may be repetitions and overlapping datasets across these repositories, resulting in a redundant and fragile ecosystem. Furthermore, each repository currently uses its own proprietary format, which is not directly interoperable with other software libraries.
As research and development in hypergraph-based methods expand across diverse domains such as data mining, computational biology, and network analysis, the absence of a common data interchange format has become an obstacle to interoperability and reproducibility. Most existing libraries and tools rely on ad hoc representations of hypergraphs, making it difficult to share data, compare results, or construct integrated workflows. In practical applications, hypergraphs typically pass through multiple stages of analysis such as construction, transformation, and visualization, with each stage potentially handled by a different tool. Without a standardized format, these transitions require custom conversion scripts and increase the risk of semantic loss or data misinterpretation.
HIF addresses this challenge by offering a consistent, expressive, and extensible JSON-based schema for encoding both the topology and metadata of hypergraphs. Its language-agnostic design enables seamless exchange between libraries implemented in Python, C++, Julia, JavaScript, and other environments, promoting modular, interoperable workflows while reducing redundant engineering effort. HIF serves as a unifying layer for the higher-order network science ecosystem, allowing software tools and pipelines to communicate through a shared representation. This facilitates reproducible research and fosters a more cohesive, collaborative development environment, similar to the role of the Open Neural Network Exchange (ONNX) format in machine learning [36] and GraphML in traditional graph processing [14].
II Higher-Order Networks
The term ”higher-order network” has become associated with a variety of data objects related to hypergraphs. The literature on hypergraphs is extensive, and there are multiple ways to define them; hypergraph representations must specify how to handle issues like the role of duplicate hyperedges, self-loops, empty hyperedges, isolated vertices, and metadata. We refer the reader to two classic publications about hypergraphs [37, 38] as well as more recent treatments of the larger space of higher-order networks [39, 40, 41]. And while the foundational mathematics of hypergraphs continues to evolve, including in a category-theoretical framework [42], our approach is to seek an appropriate general grounding sufficient to accommodate the widest range of both current hypergraph modeling approaches and current data science needs, while avoiding some of the more subtle complexities which can arise. To that end, we now introduce just a few concepts of the most direct relevance to HIF formally, while noting both similarities and differences with some of the most widely circulated concepts as used in the literature.
For our purposes, a hypergraph is a system $H=(V,E,I)$ where $V=\{v\}$ is a finite, non-empty set of vertices or nodes, $E=\{e\}$ is a finite, non-empty set of edges or hyperedges, and $I\subseteq E\times V$ is a set of incidences, that is, pairs $(v,e)$ of nodes and edges.
It may be valuable to understand how this definition may differ from some readers’ expectations. First, note that an edge $e\in E$ can be mapped to the collection of vertices with which it has an incidence: $e\rightarrow\{v\in V:(v,e)\in E\}$. With this in mind, we will also feel free to simply think of an edge as a subset of vertices, and notate $e\subseteq V$, thereby recovering the traditional sense of a hyperedge as a collection of vertices of arbitrary size. And indeed, we could also notate $H=(V,E)$ and simply understand that each edge $e\in E$ is a subset $e\subseteq V$. But what’s critical to note is that then $E$ is not actually a set, but rather, using our definition $E$ is an ”indexed family” of hyperedges possibly containing duplicates: there can be two distinct edges $e,f\in E$ which map to the same set of vertices, so that $e=f$. So while mathematically, our structure may be properly called a multi-hypergraph, for this paper, we will simply call these all hypergraphs, while still keeping this issue firmly in mind.
Rather more well understood is how any particular edge is a subset $e\subseteq V$ of arbitrary size. If $e$ maps to only two vertices, then that $e$ is a traditional (undirected) graph edge; and indeed, if this is true for all edges $e\in E$, then $H$ is called a 2-uniform hypergraph, which is just a graph: all graphs are hypergraphs. But $e$ may also have more than two nodes (the traditional sense of ”higher-order”), one node (singleton hyperedges), or even none (empty hyperedges are permitted).
Another prominent structure in higher-order network theory is that of an abstract simplicial complex (ASC). An ASC is a hypergraph, but one which is ”downward closed” in that for every hyperedge $e\in E$, then all nonempty collections of vertices $f\subseteq e$ contained in $e$ are also in the hypergraph: $f\in E$.
Essential to HIF is that it supports attributed hypergraphs: nodes $v\in V$ and edges $e\in E$ have arbitrary properties. But of great significance is that in addition, HIF also supports properties on incidences $(v,e)\in I$. In other words, a node $v\in V$ can have one property when associated with an incidence $(v,e)$ involving one edge $e\in E$ in which it is contained ($v\in e$), but could have quite a different property when associated with a different incidence $(v,f)$ for a different edge $f\in E$, for which also $v\in f$.
Incidence properties are essential for supporting critical features like ”edge-dependent weights” [43] used in, e.g., random walk models on hypergraphs [44]. But another critical use of incidence properties in HIF is to represent directed hypergraphs [45, 46]. Directed hypergraphs comprise directed hyperedges each of which is an ordered pair of sets: the ”tail”, specifying the inputs in the multi-way interaction, and the ”head”, specifying the outputs. In contrast to prior literature requiring that the head and tail be disjoint [47], here we generalize this framework and allow nodes to belong to both the tail and head. An incidence property can be used to indicate whether, for any particular hyperedge $e\in E$, a vertex $v\in e$ is in the tail, the head, or both.
III The Hypergraph Interchange Format
The Hypergraph Interchange Format is a standard for storing higher-order networks in JavaScript Object Notation (JSON) format. JSON is chosen for its interpretability; while there are more efficient formats such as Parquet, CSV, GraphQL, etc., these data structures either lack human-readability or support for the rich attributes often accompanying higher-order data. JSON is supported by all major programming languages and JSON validation libraries written in Python, R, and Julia provide a user-friendly way to verify datasets, as seen in Fig. 2, making this a suitably popular file format. HIF offers support not only for network-level metadata, but also for attributes corresponding to nodes, hyperedges, and incidences and it does so in an interpretable way. This specification also handles several different types of higher-order networks and is flexible enough to accommodate additional higher-order representations in the future.
HIF is the first cross-platform standardized file format for datasets of higher-order networks, hoping to unify the higher-order network science community around a set of standards and best practices. This standard was crystallized through exploration of the practical needs of the community and limitations of implementation. Not only do we describe a data standard, but we also provide a workflow for updating this schema as the needs of the higher-order network science community evolve and JSON schema specifications are updated.
III.1 The Hypergraph Interchange Format specification
Figure 1: The hierarchy of properties in the HIF JSON Schema.
Field Type Description
Top-level attributes
network-type (optional) enum ”directed”, ”undirected”, or ”asc”
metadata (optional) object network-level metadata
incidences (required) array lists of records representing edge-node pairs
nodes (optional) array lists of node records
edges (optional) array lists of edge records
Entries in the incidences array
edge (required) string or integer an edge ID
node (required) string or integer a node ID belonging to that edge
weight (optional) number incidence weight
direction (optional) enum ”head” or ”tail”
attrs (optional) object library-specific non-structural attributes
Entries in the nodes array
node (required) string or integer a global node ID
weight (optional) number node weight
attrs (optional) object library-specific non-structural attributes
Entries in the edges array
edge (required) string or integer a global hyperedge ID
weight (optional) number edge weight
attrs (optional) object library-specific non-structural attributes
Table 1: A description of the HIF schema. The schema is separated into five types of network data: the type of the higher-order network (encoded in network-type), network-level metadata (encoded in metadata), lists of nodes and edges with their attributes (encoded in nodes and edges, respectively), and structural information comprising information about the constituent nodes, hyperedges, and incidences.
The HIF specification organizes data about a higher-order network in five different categories via five top-level fields in the JSON file: network-type, metadata, nodes, edges, and incidences. The text corresponding to the network-type field indicates the type of the higher-order network and offers additional guidance for processing the dataset which we will discuss in detail below. The data corresponding to the metadata field is a dictionary-like object representing network-level attributes such as the name of the dataset, when it was collected, the author of the dataset, corresponding references, and other high-level information enabling greater dataset interpretability. The nodes and edges fields store data records for the nodes and edges, respectively. These records can be used to specify metadata corresponding to each node and hyperedge, or as we will discuss in more detail below, to specify empty edges or isolated nodes. Lastly, the incidences field stores data records detailing node-hyperedge relationships (known as an incidence). Below we detail more of the structure as well as some of the implications of this specification.
We start with the incidences data, because this is the only top-level attribute which explicitly encodes the complex relationships between nodes via hyperedges. The incidences field is the only required field in HIF-compliant files, and a file only including this field corresponds to a higher-order network without any network, node, or hyperedge attributes. The incidence data are composed of a list of entries known as records, where each record indicates a relationship between a node and a hyperedge, known as an incidence. For example, {”edge”: 1, ”node”: 3} indicates that node 3 belongs to hyperedge 1. The collection of all records where ”edge” is 1 forms a hyperedge, e.g., the collection of records {”edge”: 1, ”node”: 3}, {”edge”: 1, ”node”: 4},{”edge”: 1, ”node”: 7} corresponds to a hyperedge comprising the interaction between nodes 3, 4, and 7. Additional structural features which can be specified are weight and, in the case of directed hypergraphs, direction. These properties are added as top-level attributes in each record; e.g., {”edge”: 1, ”node”: 3, ”weight”: 2} specifies an incidence with an edge-dependent node weight of 2 and {”edge”: 1, ”node”: 3, ”direction”: ”tail”} indicates that node 3 is in the tail of directed hyperedge 1. Incidences can also have properties, which indicate hyperedge-dependent nodal attributes. All of these attributes are bundled as dictionary-like objects in each incidence record under the attrs field, e.g., {”edge”: 1, ”node”: 3, ”attrs”: {”role”: ”PI”}} might indicate node 3’s role as a PI on project 1.
Similarly, we can specify attributes for nodes and hyperedges by creating records corresponding to the nodes and edges fields, respectively. For example, {”node”: 0, ”attrs”: {”height”: 176, ”weight”: 143}} corresponds to node 0 with a height of 176 and weight of 143, and {”edge”: 1, ”attrs”: {”duration”: 93, ”setting”: ”coffee shop”}} corresponds to an 93 minute interaction at a coffee shop. Another use of these fields is to specify empty hyperedges and isolated nodes which don’t participate in these incidence relationships and thus are not listed in the incidence records. We can add these nodes and edges as node and edge records with or without attributes to ensure that they are present in the higher-order network.
The network-type field can take one of three values: ”asc”, ”directed”, and ”undirected”. This field offers additional guidance for processing the higher-order network. For example, using the ”asc” keyword indicates that, even if all subfaces are not specified, that the network should be downward closed. The best practice for storing a simplicial complex is to only store the maximal faces and subfaces with attributes to minimize the storage required, but this is a guideline, not a hard requirement. When using the ”directed” keyword, this indicates that the library should expect the ”direction” field indicating whether the incidence forms part of the head or tail of the directed hyperedge. When the network-type is not specified, it is assumed that the type is an undirected hypergraph.
III.2 Checking HIF compliance
A central contribution of the HIF project is a JSON schema allowing simple validation of datasets against the standard. There are simple JSON validators in several different languages and we illustrate simple ways to validate datasets against the HIF Standard in Fig. 2.
Figure 2: Example code snippets for validating the schema in different languages. In all cases, url=“https://raw.githubusercontent.com/pszufe/HIF-standard/main/schemas/hif_schema.json” and filepath is the local filepath of the file being validated.
We emphasize that the validating the schema should be the first step in the implementation of an independent software packages which utilize this standard are ultimately responsible for reading and writing datasets to and from the appropriate network representation. For example, the standard uses the ”asc” keyword to specify that every interaction present in the hypergraph is downward closed, but checking the dataset against the schema doesn’t explicitly check this. In addition, best practices are that only the maximal simplex faces and subfaces with corresponding attributes are stored for simplicial complexes, but this informal standard is not enforced by the schema. Lastly, while end users are able to, in principle, specify the direction keyword for network datasets even if the network-type keyword is not directed and to omit the direction keyword even when network-type is directed, for simplicity, we simply say that direction field is optional.
III.3 Updating the schema
Changes made to the schema are documented in a CHANGELOG.md document. When a new version is released, the version process is followed:
  1. 1.
    Decide on the new version number based on the changes made since the last version and the Semantic Versioning guidelines [48].
  2. 2.
    Copy the schema in the hif_ schema.json file to a file named hif_ schema_ <version>.json where version indicates the new version.
  3. 3.
    Add the changes made since the last release to CHANGELOG.md in a section with the new version as the name.
  4. 4.
    Upload the new stable version to Zenodo as a persistent reference.
The hif_ schema.json schema will always have version ”latest” and all unit tests are based on this schema.
IV Integration with software libraries
A strength of the HIF standard is its integration with five prominent software libraries for higher-order network analysis: the Hypergraph Analysis Toolbox (HAT) [24], Hypergraphx [26], HyperNetX [27], SimpleHypergraphs.jl [29], and XGI [31]. Here we detail the ways in which each library supports the HIF standard as well as brief descriptions of their use.
IV.1 Hypergraphx
Hypergraphx (HGX) is an open-source Python library designed for the analysis of complex systems characterized by higher-order interactions. HGX offers an extensive collection of tools and algorithms for constructing, manipulating, analyzing, and visualizing hypergraphs, enabling users to capture the rich structure and study the dynamics of real-world systems beyond pairwise relationships. HGX enables seamless storage and conversion of higher-order data across multiple representations, including hypergraphs, simplicial complexes, bipartite networks, line graphs and clique expansions, all while supporting hyperedges with weights, direction, sign, temporal dynamics and multiplex layers. The library provides tools for basic statistical characterizations (e.g., hyperedge size distribution), higher-order centrality, motif analysis (with scalable sampling), community detection, significant interaction filtering, and synthetic hypergraph generation. HGX also supports the simulation of dynamical processes and offers advanced visualization functionalities for exploring the structure of real-world systems.
On the data and I/O side, HGX is accompanied by HGX-data, a collection of ready-to-use datasets for higher-order network analysis in Python, including collaboration, face-to-face, biological data, and more. HGX offers dedicated input and output functions, including a binary format to enable fast and efficient I/O over large-scale higher-order datasets. Hypergraphx has integrated HIF support through the hypergraphx.readwrite.read_ hif function, allowing users to load hypergraphs directly from HIF-compliant JSON files and automatically handling the conversion to HGX’s internal data structures. Similarly, the hypergraphx.readwrite.write_ hif function enables users to export an HGX Hypergraph object to the HIF format, preserving both the structural information and any associated node, edge, or incidence attributes. This interoperability allows users to seamlessly leverage HGX functionalities while moving back and forth across different libraries. In the future, HIF will accommodate additional hypergraph representations available in Hypergraphx, including multiplex hypergraphs. Moreover, HGX-data is planned to be fully HIF-compatible in order to be accessed by the whole higher-order software ecosystem.
An example demonstrating how to load a hypergraph from a HIF file, leverage HGX functionalities such as motif analysis, and export the data is provided in Fig. 3.
Figure 3: Example demonstrating how to load a HIF-compliant file, use Hypergraphx (HGX) functionalities and export the file.
IV.2 HyperNetX
HyperNetX (HNX) is a python package focused on developing tools for the analysis and visualization of hypergraphs. HNX provides a user-friendly interface for creating, manipulating, and studying hypergraphs, with an emphasis on visualization capabilities that capture higher-order relationships. Some of the key features of HNX include:
  • A suite of analytical functions, including degree sequences, connectivity measures, and centrality metrics.
  • Tutorials for both basic and advanced hypergraph analysis methods with a contributors guide for the addition of new modules and tutorials.
  • Visualization tools for rendering hypergraphs in various layouts, highlighting different aspects of their structure.
  • Support for incidence, node, and hyperedge attributes, allowing for richer data representation.
HyperNetX has integrated HIF through the hnx.read_ hif and hnx.to_ hif functions: hnx.read_ hif loads hypergraphs directly from HIF-compliant JSON files, handling the conversion to HNX’s internal data structure and hnx.to_ hif exports an HNX hypergraph object to the HIF format, preserving both the structural information and any associated node, edge, or incidence attributes. An example demonstrating how to load and save an HNX hypergraph is shown in Fig. 4 where the filename is provided. Alternatively, the JSON object can be loaded first and then the JSON object converted to an HNX hypergraph can be done.
Figure 4: Example demonstrating how to load and export and HIF file using HyperNetX (HNX).
IV.3 SimpleHypergraphs.jl
SimpleHypergraphs.jl is a Julia package designed to efficiently work with hypergraphs [29]. Its key capabilities include (1) creating hypergraphs from incidence matrices, random hypergraph models, a Graphs.jl graph, or external data; (2) manipulating hypergraphs, (3) analyzing hypergraphs with fast algorithms for computing, for example, connected components or modularity based clustering; (4) visualizing hypergraphs; (5) integration with the Graphs.jl Julia package via bipartite graphs or a hypergraph projection; and (6) and file I/O. SimpleHypergraphs.jl is designed to support (1) two-way of storage of node and hyperedge incidence information, which enables efficient traversal of the hypergraph in both directions — from nodes to hyperedges and vice versa; (2) hyperedge-dependent node weights, i.e., incidence weights; (3) node and hyperedge metadata; (4) compatibility with Julia’s AbstractMatrix interface, enabling efficient manipulation of hypergraphs using standard Julia matrix operations; and (5) lazy representation of hypergraph views which are compatible with the AbstractGraph interface of the Graphs.jl library, enabling seamless integration and direct usage of Graphs.jl algorithms. Figure 5 presents an example of SimpleHypergraphs.jl library usage with HIF.
Figure 5: The example code demonstrates creating a hypergraph with node and hyperedge metadata represented as strings and exporting the hypergraph to a HIF-compliant JSON file.
IV.4 Hypergraph Analysis Toolbox
The Hypergraph Analysis Toolbox (HAT) [24] is a general-purpose software for constructing, visualizing, and analyzing hypergraphs and higher-order structures, with a focus on their structure and dynamics. HAT implements tensor based methods for the analysis of higher-order observability/controllability [49, 50], coupling [51], and structural properties suitable for analysis of a wide range of data [52, 53]. It provides HIF-compliant read and write capabilities in the HAT.Hypergraphs module. Figure 6 presents a sample using the HAT library. The Hypergraph.to_ hif method converts the hypergraph object back to a Python dictionary according to the HIF schema, which can be saved to a JSON file.
Figure 6: The example code demonstrates creating a HAT hypergraph from an example HIF JSON file and exporting it back to Python.
IV.5 XGI
The CompleX Group Interactions (XGI) [31] software package provides support for hypergraphs, directed hypergraphs, and simplicial complexes. It provides a comprehensive suite of analysis and visualization tools, including algorithms, generative higher-order network models, linear algebra representations, conversions between many different data structures, and an integrated statistical interface. Integrated with this software environment is the XGI-DATA repository, which hosts many higher-order datasets. XGI provides not only visualization and analysis tools for higher-order networks but the ability to read and write numerous different file formats such as bipartite edge lists, hyperedge lists, and incidence matrices.
XGI currently supports the HIF standard in two ways: first, it provides HIF-compliant read and write capabilities and second, it provides HIF-compliant datasets in the XGI-DATA collection. XGI has implemented an hif sub-module in its readwrite module, providing two methods: read_ hif and write_ hif. These methods allow XGI to store undirected hypergraphs, directed hypergraphs, and simplicial complexes according to the HIF standard. When storing directed hypergraphs, XGI makes use of the direction field in the incidence records. When storing simplicial complexes, XGI only stores the maximal faces and hyperedges with associated attributes, with the expectation that the asc keyword will indicate that XGI should enforce downward closure when reading in the file. The HIF standard is robust to the future development of XGI; for example, it implements features not yet supported by XGI. One example is edge-dependent node attributes, also known as incidence attributes, which capture properties of physical systems such as stoichiometric coefficients in chemical reactions and interaction-dependent social roles.
Figure 7: The code example demonstrates using XGI to read and write HIF-compliant files. First, a hypergraph with hyperedges $\{1,2,3\}$, $\{2,3,4\}$, and $\{1,4\}$ is generated and written to an HIF-compliant JSON file. Second, the hypergraph is read from that file, casting the names of the nodes from integers to strings, and the hyperedge list is returned.
The load_ xgi_ data function in XGI supports reading HIF-compliant files from the XGI-DATA repository and several datasets in XGI-DATA are stored in XGI-DATA. Future plans include converting all datasets to HIF, to make XGI-DATA a shared resource for all higher-order software libraries.
V Case study
To illustrate the practical application and interoperability enabled by the Hypergraph Interchange Format (HIF), we present a case study. Figure 9 provides an overview of how the HIF standard facilitates the exchange of hypergraph data across several hypergraph analysis software packages — Hypergraphx, HyperNetX, SimpleHypergraphs.jl, and XGI — enabling more integrated and sophisticated data analysis pipelines.
V.1 Dataset
We create the publications.hif.json HIF-compliant file and interact with this dataset by loading the file through each software library’s API.
Figure 8: Illustration of the full publications hypergraph including disconnected components.
We use a publication dataset for our case study because of its rich attributes and tractable size. This dataset consists of open-source publications with the keyword ”Hypergraph” and was collected from ArXiv, BioRxiv, the DBLP computer science bibliography, and the U.S. Office of Scientific and Technical Information (OSTI). The resulting hypergraph is composed of hyperedges, which represent scientific publications, and nodes, which represent authors. Each hyperedge (publication) has a number of attributes including funding agencies, abstract, publication date, keyword tags, and the source, while each node (author) has an associated institution have the attributes of institutions (There are no incidence attributes). The hypergraph has 1,960 nodes (authors) and 533 edges (publications). The following analysis of this data using each of the hypergraph packages is provided in the notebook HIF-demo.ipynb in the tutorials folder of the HIF GitHub page  [^2] See https://github.com/pszufe/HIF-standard.. We briefly describe examples of the types of analysis that one can perform on this dataset for each software package.
Figure 9: The Hypergraph Interchange Format enables interoperability between many popular higher-order network analysis libraries, as seen in this case study which analyzes a publication dataset. The notebook that creates each visualization can be found in the HIF-demo.ipynb file. In the center, we display the HIF logo. In panel (a), we use HAT to compute the nonlinear eigenvector centralities of the nodes and hyperedges. In panel (b), we use Hypergraphx to visualize the largest connected component of the hypergraph (i), with node colors indicating the community structure. We also visualize the most frequent patterns of group interactions involving three (ii) and four nodes (iii). In panel (c), we use HyperNetX to solely examine the largest component of the hypergraph, compute the closeness centrality for each hyperedge, and color each hyperedge according to its centrality as we can see from the color bar. In panel (d), we use SimpleHypergraphs.jl to find the largest component, compute modularity-based community detection, and to visualize the pairwise network projected from the hypergraph, where each hyperedge becomes a clique. In panel (e), we use XGI to compute nodal statistics using the statistics interface; in the visualization, the node size is scaled by the nodal degree and the node color corresponds to the clique eigenvector centrality (CEC) [55]. Similarly, in the top inset figure, we see the degree (solid line) and average neighbor degree (dashed line), while in the bottom inset figure, we see the CEC (solid line) and the H-eigenvector centrality (HEC) [55, 56] (dashed line). Note that the node labels are sorted in descending order according to the degree (top) and CEC (bottom).
V.1.1 Hypergraph Analysis Toolbox
For HAT we demonstrate the usefulness of analyzing the publications hypergraph using the eigenvector centrality measure. This allows us to analyze node and edge centrality. In 9(a), we compute the centralities of each node and edge.
V.1.2 Hypergraphx
To demonstrate the capabilities of Hypergraphx, we consider our example publications dataset and perform an exploratory analysis combining community detection, motifs, and visualization. Specifically, in Fig. 9(b), we apply (overlapping) community detection to the largest connected component, identifying groups of closely related nodes [57], and visualize the structure of the resulting hypergraph. We further use Hypergraphx to mine the frequency of subhypergraph patterns of sizes 3 and 4 [58], describing the hypergraph of publications at its microscale and providing insights into the structural building blocks of higher-order interactions present in the data [59]. This use case illustrates how Hypergraphx can serve as a key component of an integrated workflow within the higher-order network ecosystem, enabling complex tasks such as motif analysis and community detection to be performed efficiently and in combination with other tools.
V.1.3 HyperNetX
For HyperNetX, we demonstrate the usefulness of analyzing the main component of the publications hypergraph using the $s$-closeness centrality measure for $s=1$, visualized as an Euler diagram in Fig. 9(c). We did not use the full hypergraph for this centrality analysis as it is too dense and disconnected to visually discern hyperedge centrality. The $s$-closeness centrality measure shows the closeness of each hyperedge (in this case, a publication) based on the shared authors. The color scale indicates the centrality value, with warmer colors (yellow) representing hyperedges that are more central, meaning they share more authors with other publications. This analysis can highlight influential publications that are central to the collaborative structure of this publication hypergraph.
V.1.4 SimpleHypergraphs.jl
For the analysis with SimpleHypergraphs.jl we have selected the largest connected component of the citation hypergraph that has $108$ hypervertices and $30$ hyperedges. Next, we use library’s functionality hypergraph modularity-based clustering introduced in Ref. [60] to identify clusters within the citation hypergraph. Subsequently, we project the hypergraph to a pairwise projection where every hyperedge becomes a clique. We visualize the graph in Fig. 9(d), scaling node sizes by their degrees and using colors to denote community labels. All communities with a single node are colored in black.
V.1.5 XGI
A strength of the XGI software package is its statistics interface, allowing practitioners to easily visualize, output as common Python data types, and compute statistics of node and edge properties. In Fig. 9(e), we see a hypergraph with its nodes colored and sized according to its degree and the average degree of its neighbors. XGI can also easily filter datasets by hyperedge size according to the framework in Ref. [61]. Another strength of the XGI software package is its ability to quickly remove data artifacts with its cleanup method which has the ability to remove singletons, isolated nodes, multi-edges, and disconnected components. In Fig. 9(e), we see the largest component of the hypergraph without any multi-edges, isolated nodes, or singletons.
VI Discussion
In this paper, we introduced the Hypergraph Interchange Format as a standard for sharing hypergraph data across different scientific workflows. We demonstrated its effectiveness with a case study showcasing the use of several software packages to analyze a single dataset in a variety of ways, each catering to a different library’s strengths. In the current fragmented open-source ecosystem, the HIF standard offers interoperability between many higher-order network analysis libraries.
As part of this effort, we surveyed the landscape of open-source software for existing standards and potential adopters, reaching out to more than 20 open-source projects with support for hypergraphs and simplicial complexes. Our exploration validated the hypothesis that the community of higher-order networks needs an open standard and that the HIF design supports many use cases. Future work will include integration with more software libraries to improve its reach in the higher-order network science community.
The HIF standard currently emphasizes human readability and support for richly attributed data, but one can imagine more efficient ways for storing datasets which could be implemented in future iterations of HIF. For example, Javascript Object Notation Lines (JSONL) is a specification defining how to transmit a top-level list of JSON objects as a series of one-line JSON documents, enabling streaming file input/output (I/O). When solely considering performance, one could consider a database architecture with a query language such as GraphQL as a performant alternative for storing and retrieving higher-order network data. As large higher-order datasets become increasingly more available, efficient storage and handling of higher-order network data will become more and more necessary.
Lastly, while the HIF standard directly supports undirected and directed hypergraphs and simplicial complexes, support for additional higher-order representations will make HIF more widely useful. Examples include hypergraphs with ordered and partially ordered sets (posets) which are useful in legal analysis [62] and the Science of Science. Several types of higher-order networks are implicitly supported by the HIF standard: signed hypergraphs which assign a sign +/- to each hyperedge; multilayer hypergraphs which assign nodes and hyperedges to different layers representing, for example, modes of transportation, relationship types, or social media platforms; and temporal hypergraphs where nodes and hyperedges have associated time intervals in which they are active. The specific structural information necessary for each of these representations can be stored in the attrs field, but the schema excludes these as expected JSON fields. In the future, adding explicit support for these fields will improve compatibility with libraries such as ASH [23], Raphtory [63], and Reticula [28]. In addition, we recommend that when releasing new higher-order datasets that an HIF-compliant file is included in the release (for example, see Ref. [64]).
While there are many avenues for future growth of HIF, we believe that this is a significant first step in the standardization of the higher-order network data sharing ecosystem. It has a direct impact on the reproducibility of published works that adopt the standard. HIF will facilitate seamless data exchange, cross-software collaboration, and more compact higher-order data analysis pipelines and will be a significant advance for the coming years of higher-order network science.
Data Availability
All datasets, resources, documentation, and schemas are available on GitHub footnotemark: . The stable schema is available on Zenodo [65] and subsequent stable versions will be released on Zenodo.
Contribution statement.
Author order is alphabetical. N.W.L. led the project. All authors designed the HIF standard and wrote the manuscript.
Funding.
N.W.L. acknowledges support from the University of Virginia Prominence-to-Preeminence (P2PE) STEM Targeted Initiatives Fund, SIF176A Contagion Science.
This work acknowledges support from Pacific Northwest National Laboratory with information release number PNNL-SA-213514